OpenStudy (anonymous):

how to evaluate sin215 ?

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

i didnt get that, sorry

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

sin (180+45)= sin180 +sin45 = 0 + (1/sqrt 2) = 1/sqrt 2

OpenStudy (anonymous):